OpenStudy (kainui):

I found an interesting problem: ABCD x E = DCBA so each letter is really a digit, and they're all different digits. So it could be ABCD=1234 or something like that. Any ideas on how to solve this thing?

OpenStudy (anonymous):

What we have to find

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

What are the numbers that make this a true statement is all.

OpenStudy (anonymous):

so if A is 2 then B should be 3 for example

OpenStudy (anonymous):

?

OpenStudy (kainui):

Here's my attempt, it's obviously wrong but fun to look at. \[(A*10^3+B*10^2+C*10+D)*E=D*10^3+C*10^2+B*10+A\]\[AE*10^3+BE*10^2+CE*10+DE=D*10^3+C*10^2+B*10+A\] Collect all the terms with the same power of 10 (obviously wrong but follow along) AE=D, BE=C, CE=B, and DE=A take the log of all these equations to get: logA+logE-logD=0, logB+logE-logC=0, etc... then plug them into a matrix: \[\left[\begin{matrix}1 & 0&0&1&-1 \\ 0 & 1&-1&0&1 \\0 & -1&1&0&1 \\-1 & 0&0&1&1 \end{matrix}\right]\left(\begin{matrix}logA \\ logB\\logC\\logD\\logE\end{matrix}\right)=\left(\begin{matrix}0 \\ 0\\0\\0\end{matrix}\right)\] row reduce the augmented matrix and you solve to get A=D, B=C, and E=1. So obviously this is wrong because we can't have any of the digits equal to each other, but it works if that wasn't a constraint. ABBA*1=ABBA So that's the best I've got lol.

OpenStudy (kainui):

A, B, C, D, and E are all just digits between 1 and 9 and none of them are equal. They can be anything as long as they make the equation work out. So for instance, if A=2 and B=4 and C=5 then AB x C = 24*5=96

ganeshie8 (ganeshie8):

i think that works only if there was no carry any where in the multiplication by E

OpenStudy (kainui):

Yeah the problem is the carrying. So I could either change base or try to figure out how to use the modulus function maybe? Idk. DEmod10=A CEmod100=B etc... I think that's how it works lol idk.

OpenStudy (asnaseer):

Assuming A, B, C, D and E are distinct digits, then we can say that: E > 1 and therefore that: A < 5

OpenStudy (anonymous):

what do u mean by E ?

OpenStudy (asnaseer):

if A=1 then D*E must be a product that ends in a "1", the only combinations are: 7x3 OR 3x7 which means we either have: 1BC7 x 3 = 7CB1 (invalid combination) or: 1BC3 x 7 = 3CB1 (cannot be possible) therefore A > 1 and therefore E < 5 so now we know that: 1 < A < 5 1 < E < 5

OpenStudy (asnaseer):

now we need to find combinations of DxE that yield A as their units digit A=2 => DxE = 4x3 OR 3x4 OR 4x8 OR 8x4 A=3 => DxE = no valid combinations A=4 => DxE = 2x7 OR 7x2 OR 3x8 OR 8x3

OpenStudy (kainui):

E is just another number. An example of a wrong solution would be: 2583*7=3852.

OpenStudy (asnaseer):

which narrows A down to either 2 OR 4

OpenStudy (asnaseer):

Working through this reduced set I get one valid solution: 2178 x 4 = 8712

OpenStudy (kainui):

I don't see why A can't be 1, 2, 3, or 4.

OpenStudy (asnaseer):

do you agree that if A=1 then D*E must be a product that ends in a "1"?

OpenStudy (anonymous):

(D+10C+100B+1000A)E=A+10B+100C+1000D D*E=A C*E=B B*E=C A*E=D

OpenStudy (anonymous):

so it work if :- A/D=B/C i got a ratio :3 lets make sure of it (so we must have no zero )

OpenStudy (anonymous):

ohh we should know E :-|

OpenStudy (asnaseer):

@BSwan - you are assuming there will be no carry

OpenStudy (kainui):

Here are some equations I noticed: DE mod10 = A DE mod100 - DE mod10 + CE mod10 = B CE mod100 - CE mod10 + BE mod10 = C BE mod100 - BE mod10 + AE mod10 = D AE mod100 - AE mod10 = 0 These all come from making sure each digit lines up. Noticing a couple extra things, since we don't have repeating digits, E=/=1 For any situation the highest possible combination of digits can be 9*8, which is 72. So this means that if X and Y are any digits, then XY mod100 - XY mod10 <= 7 I combined all first equations at the top and got: A + B + C + D = AE mod100 + BE mod100 + CE mod100 + DE mod100 then using that last inequality I came up with: A + B + C + D <= 4*7 + AE mod10 + BE mod10 + CE mod10 + DE mod10 From here I have absolutely no clue if this is in the right direction or not because I really don't know how to use modulus with algebra.